According to insurance records, a car with a certain protection system will be recovered 88% of the time. If 500 stolen cars are randomly selected, what is the mean and standard deviation of the number of cars recovered after being stolen?
Answer by Boreal(15235)

mean is n*p=500*0.88=440 cars
variance is np(1-p)=440*0.15=66 cars^2
sd is sqrt(V)=8.12 cars